OpenStudy (anonymous):

You're driving down the highway late one night at 18m/s when a deer steps onto the road 39m in front of you. Your reaction time before stepping on the brakes is 0.50s , and the maximum deceleration of your car is 12m/s2 . How much distance is between you and the deer when you come to a stop? =17m What is the maximum speed you could have and still not hit the deer? = ?

OpenStudy (anonymous):

You have to find the maximum initial velocity here, when with a maximum deceleration of 12 m/s^2 you can stop the car just before the deer. Also, you found the distance, s=17 m. So, enter all the value into the very common eqn of kinematics- v^2=u^2-2*a*s u=initial velocity v=0=final velocity

OpenStudy (anonymous):

Wait, I made a mistake, Let me think a while.

OpenStudy (anonymous):

I think you made a wrong calculation. The first answer should be= 16.5 m

OpenStudy (anonymous):

Try s=30 m in the above eqn. Because the brak

OpenStudy (anonymous):

*Because the brake was being applied during this time.

OpenStudy (loser66):

@idban , he is right, he takes significant figure (2 sg.f)

OpenStudy (loser66):